# Personalized Federated Learning under Model Dissimilarity Constraints

### v1

@@ Abstract
One of the defining challenges in federated learning is that of statistical heterogeneity among clients. We address this problem with Karula, a regularized strategy for personalized federated learning, which constrains the pairwise model dissimilarities between clients based on the difference in their distributions, as measured by a surrogate for the 1-Wasserstein distance adapted for the federated setting. This allows the strategy to adapt to highly complex interrelations between clients, that e.g., clustered approaches fail to capture. We propose an inexact projected stochastic gradient algorithm to solve the constrained problem that the strategy defines, and show theoretically that it converges with smooth, possibly non-convex losses to a neighborhood of a stationary point with rate $O(1/K)$. We demonstrate the effectiveness of Karula on synthetic and real federated data sets.
@@ 1 Introduction
Federated learning (FL) is a learning paradigm in which models are trained on data owned by clients in a network which is connected to a central server. As a response to the increasing presence of private and non-centrally stored data, e.g., data stored on mobile devices or hospital databases, the goal of federated learning is to learn models in a collaborative and communication-efficient manner, without exchanging client data. Federated averaging (McMahan et al.,, 2017) is an example of a FL algorithm, in which clients locally perform gradient steps on a global model, and send updated model parameters to the server, which then averages and returns the updated global model to the clients. There are many examples where FL has been applied successfully, leading to greater communication-efficiency and data privacy, e.g., in mobile keyboard prediciton (Hard et al.,, 2018), autonomous driving (Nguyen et al.,, 2022), and healthcare (Qayyum et al.,, 2022). However, FL faces several challenges that arise due to client hetereogeneity, be it in terms of models, hardware, or data. In the case of data heterogeneity, the client data follow different statistical distributions and may have varying number of observations. This means that the performance of standard FL systems, which learn a single global model, can suffer significantly, both in terms of convergence and generalization. At the same time, due to limited data, non-collaboratively training separate models also leads to poor generalization.
@@
Therefore, in settings where data heterogeneity is significant, federated learning systems should ideally learn personalized models in a collaborative manner where each model borrows strength from somewhere else. Personalized federated learning (PFL) strategies attempt to do just that. We will divide these strategies into two categories: (1) strategies that learn local models that borrow strength from a global model; (2) strategies that learn local models that borrow strength from clients identified as similar at the server-level. We refer to the two categories as globally anchored personalization and peer-based personalization, respectively. Globally anchored strategies are appropriate in applications where we expect each client to have limited dissimilarity with any other client in the network, such as in next-word prediction on mobile devices. However, in applications such as recommender systems, distributional differences between two given clients can be very large. Peer-based methods address this issue by attempting to only learn from clients that are similar. One approach to peer-based personalization is to cluster clients, prior to training or concurrently. The cluster model of statistical heterogeneity does however fail to capture more complex overlaps between clients. As an example, in a music recommendation application, suppose one user (client) is only interested in jazz and classical, another in jazz and funk, and yet another in funk and rock. Clearly, no hard clustering of these users will capture these relationships accurately. Other approaches to peer-based personalization allow for more complex relationships between the local models of clients, but to the best of our knowledge no prior work incorporates distributional similarity measures in a principled manner.
@@ 1.1 Contributions
The main contributions of this work are three-fold, and are as follows:
  • We propose a peer-based PFL strategy which learns personalized models that are centerally constrained to have limited pairwise dissimilarity. The strategy produces the limits on the model dissimilarities in a principled and completely data-driven manner using approximations of the Wasserstein distances between the empirical client distributions, motivated by forthcoming theory presented in this work.
  • We propose an efficient inexact projected stochastic gradient method for the sceme is presented, which employs variance reduction to deal with partial client participation in each communication round. Due to this variance reduction, we are able to show convergence for smooth, possibly non-convex loss at rate $O(1/K)$ to an $\epsilon$-stationary set of personalized models, where the error $\epsilon$ arises only due to stochastic gradients from the clients and inexact projection at the server level.
  • We conduct numerical studies on synthetic and real heterogeneous data sets with regression and classification tasks, and compare our strategy with a number of other FL and PFL strategies.
@@ 1.2 Outline
In §2 we describe some of the related work, relating our contribution to existing literature. Then, in §3, we introduce or proposed strategy for personalized federated learning, which is followed by a convergence analysis in §4. We present numerical experiments validating our proposed method in §5. We conclude by summarizing and discussing our findings in §6.
@@ 2 Related work
Statistical heterogeneity presents significant challenges to both traditional and distributed learning systems, affecting inference and generalization. It arises in various applications where data can be assumed to be drawn from distinct distributions. For example, in meta-analysis, studies with participants from different populations are aggregated (Melsen et al.,, 2014); in medicine, different clinical features in patients affect individual health outcomes (Machado et al.,, 2011); and in finance, markets adopt different regimes over time (Ang and Timmermann,, 2012). In the context of FL, the statistical heterogeneity among clients is generally considered one of the defining aspects of the field. The works of Zhao et al., (2018) and Karimireddy et al., (2020) show that both convergence and generalization of conventional FL algorithms are severely affected by statistical heterogeneity. The heterogeneity introduces difficulties in optimizing the global model, as updates from different clients may conflict, reducing the effectiveness of aggregation methods and slowing down convergence. The former work additionally shows that the 1-Wasserstein distance between client distributions can be used to quantify this drift.
One approach to address the challenges caused by statistical heterogeneity is stratification, whereby different models are fit for every value of some categorical parameter, which are assumed to correspond to the different underlying distributions. Stratified models allow each subgroup to be modeled independently, capturing the unique characteristics of each distribution and reducing the bias that arise from pooling heterogeneous data. In clinical trials, it is common to do this by grouping participants into different strata based on relevant clinical features, and fitting a model for each strata (Kernan et al.,, 1999). Stratification does however reduce the number of observations for each model, meaning viability of naive stratification is limited in many settings. PFL models, which adapt to individual clients rather than enforce a single global model, are a special case of stratified models where the stratification is performed at the level of individual clients.
@@
A more flexible personalization strategy is to identify clients that are similar and borrow strength from them directly rather than from a single aggregated model. One way to do this is to cluster the clients. The PFL method proposed by Ghosh et al., (2020) does this iteratively by alternating between optimizing the cluster models and clustering the clients, which is done by finding the cluster whose model achieves the smallest local loss. Regularized methods for peer-based PFL include convex clustering via a group-lasso penalty on the differences of models (Armacki et al.,, 2022) and Laplacian regularization (Dinh et al.,, 2022; Hashemi et al.,, 2024). The empirical results of Dinh et al., show that in fact Laplacian regularization with randomized weights outperform many globally anchored personalization strategies in several tasks.
@@ 3 Proposed method
In this section, we introduce Karula [^1] The name is a portmanteau of Kantorovich, Rubenstein and Laplace., a strategy for federated learning of personalized models where model differences are constrained by the distributional similarity of client data. We also propose an efficient projected stochastic gradient method with variance reduction to handle partial client participation during communication rounds. This method is also suitable for use in other FL strategies.
@@ 3.1 Main idea
Consider the FL setting with $N$ clients in a network that can communicate with a central server. To each client $i\in\{1,\dots,n\}$, we associate an underlying distribution $\mu_{i}\colon 2^{\mathcal{X}}\to{\mbox{\bf R}}_{+}$, with $\mathcal{X}$ being the feature space, and a data set $\mathcal{D}_{i}$ of $N_{i}$ independent samples drawn from this distribution. Ideally, we would learn models $\theta_{i}\in\Theta$ that minimize $f_{i}(\theta_{i})={\mathbbmss{E}}_{\xi\sim\mu_{i}}[\ell(\theta_{i},\xi)]$, where $\ell\colon\Theta\times\mathcal{X}\to{\mbox{\bf R}}$ is a loss function, and $\Theta$ is the parameter space. However, due to limited available data on each client, attempting to directly minimize $f_{i}(\theta_{i})$ leads to poor generalization, so our strategy regularizes the models by approximately solving the constrained problem
@@
Here $t\geq 0$ is a regularization parameter, each $\alpha_{i}$ is a positive weight, and each $D_{ij}\geq 0$ is a dissimilarity parameter which quantifies the dissimilarity between the $i$th and $j$th client, and are calculated before solving the minimization problem. This formulation generalizes the standard FL formulation, since by selecting $t=0$ the problem is equivalent to choosing a single global model that minimizes the average loss. In the other extreme, when $t\to\infty$, we obtain completely local models except for clients that are deemed to have identical distributions, i.e., when $D_{ij}=0$. In the sequel, we show how to choose the similarity parameters in a completely data-driven manner, based on a surrogate for the Wasserstein distance.
@@ 3.2 Model dissimilarity and Wasserstein distance
We now introduce the Wasserstein distance, and connect it to the dissimilarities between the ideal models. Suppose the data space $\mathcal{X}$ is equipped with a metric $d_{\mathcal{X}}$. Then the 1-Wasserstein distance (also known as the earth mover distance and Kantorovich–Rubenstein distance) between two distributions $\mu_{i}$ and $\mu_{j}$ is defined as
@@
where $\Pi(\mu_{i},\mu_{j})$ is the set of all probability measures with marginals $\mu_{i}$ and $\mu_{j}$. This distance is indeed a true metric on the space of distributions on $\mathcal{X}$, meaning it is symmetric, positive definite, and satisfies the triangle inequality (Villani,, 2008).
@@
The latter assumption holds naturally for some loss functions, but for others, e.g. linear regression with quadratic loss, we would need to require $\Theta$ to be bounded.
@@
A straight-forward approach to shrink the model dissimilarities would be to learn completely local models first, calculate their pairwise distances, and use them as the dissimilarity parameters. So why use a less direct approach with Wasserstein distances? In the FL setting, it is more common than not to have more features than observations on each client. In, e.g., a linear regression task with $d$ important features and clients with $N$ observations each, it is well known that the the minimax squared estimation error is $O(d/N)$ (Bickel et al.,, 2009; Raskutti et al.,, 2011). Moreover, if $N<d$, then without additional assumptions, no meaningful upper bound on the error exists. With models that have more parameters than there are features, this problem is likely to be further exarcerbated. The Wasserstein distance between two empirical distributions, however, has a rate of $O(N^{-1/d})$ (Wang et al.,, 2021, Example 1). This rate is very slow when increasing $N$, but clearly performs much better than $O(d/N)$ when $N$ is on a similar scale as $d$, or is smaller. (In Appendix TODO, we also show empirically that directly estimating the dissimilarities underperforms our proposed approach.)
@@
In a supervised learning task, a simple choice of data metric $d_{\mathcal{X}}$ is to concatenate the features and the response, and use, e.g., the Euclidean norm. Another approach is given by Elhussein and Gursoy, (2024), who propose using cosine-similarity on the features and the Hellinger distance on the response. They argue cosine-similarity is more informative in high-dimensional feature spaces due to the concentration of measures phenomenon, and that the Hellinger distance compares densities within their statistical manifold, rather than within Euclidean space.
@@
With the present motivation in mind, we can choose the dissimilarity parameters to be $W_{1}(\hat{\mu}_{i},\hat{\mu}_{j})$, with $\hat{\mu}_{i}$ being the empirical distribution of the $i$th client. Rakotomamonjy et al., (2023) present an iterative method to approximate the Wasserstein distances in a federated manner, which involes computing an optimal transport plan every iteration. This means that in our setting we would have to compute $n(n-1)/2$ optimal transport plans for a number of iterations, which is expensive both in terms of computation and of communication. We therefore propose to approximate the Wasserstein distances via linear embedding of the data sets, following works of e.g. Kolouri et al., (2020) and Liu et al., (2025).
We introduce a reference data set $\mathcal{D}_{0}=\{\xi_{1}^{0},\dots,\xi_{N_{0}}^{0}\}$ of $N_{0}$ independent samples drawn from a reference distribution $\mu_{0}$. The optimal transport plan between the empirical distributions $\hat{\mu}_{0}$ and $\hat{\mu}_{i}$ is then given by
@@
By approximating the Monge map via $M_{i}=N_{0}\pi_{i}^{\star}\mathcal{D}_{i}$, we can define an embedding $\Phi$ by $\Phi(\mathcal{D}_{i})=(M_{i}-\mathcal{D}_{0})/\sqrt{N_{0}}$ that has the property that $\|\Phi(\mathcal{D}_{i})-\Phi(\mathcal{D}_{j})\|_{1}$ defines a metric on the space of data sets, and indeed approximates the Wasserstein distance $W_{1}(\hat{\mu}_{i},\hat{\mu}_{j})$. Thus, we choose $D_{ij}=\|\Phi(\mathcal{D}_{i})-\Phi(\mathcal{D}_{j})\|_{1}$ as the dissimilarity parameters. By doing this, each client only has to compute a single optimal transport plan. Also, on the communcation side, only a reference data set and the embeddings need to be transmitted over the network once, totalling $O(N_{0}dn)$ communication.

### v2

@@ Abstract
One of the defining challenges in federated learning is that of statistical heterogeneity among clients. We address this problem with Karula, a regularized strategy for personalized federated learning that constrains the pairwise model dissimilarities between clients based on the difference in their distributions. Model similarity is measured using a surrogate for the 1-Wasserstein distance adapted for the federated setting, allowing the strategy to adapt to highly complex interrelations between clients that, e.g., clustered approaches fail to capture. We propose an inexact projected stochastic gradient algorithm to solve the resulting constrained optimization problem, and show theoretically that it converges as $O(1/K)$ to the neighborhood of a stationary point of a smooth, possibly non-convex, loss. We demonstrate the effectiveness of Karula on synthetic and real federated data sets.
@@ 1 Introduction
Federated learning (FL) is a paradigm for training models on decentralized data that remains with clients in a network coordinated by a central server. Developed in response to the growing prevalence of private, decentralized data sources—such as mobile devices and hospital databases—FL enables collaborative and communication-efficient model training without requiring the exchange of client data. One example of an FL algorithm is Federated averaging (McMahan et al.,, 2017) where clients perform local gradient steps on a global model and send updated model parameters to the server. The server then averages these and returns the updated global model to the clients. There are many examples where FL has been applied successfully, leading to greater communication-efficiency and data privacy, e.g., in mobile keyboard prediction (Hard et al.,, 2018), autonomous driving (Nguyen et al.,, 2022), and healthcare (Qayyum et al.,, 2022). However, FL faces several challenges that arise due to client heterogeneity, be it in terms of models, hardware, or data. In the case of data heterogeneity, the client data follow different statistical distributions and may have varying number of observations. This means that the performance of standard FL systems, which learn a single global model, can suffer significantly in terms of both convergence and generalization. At the same time, due to limited data, non-collaboratively training separate models also leads to poor generalization.
@@
Therefore, in settings where data heterogeneity is significant, federated learning systems should ideally learn personalized models that borrow strength from others while respecting local nuances in their data. Personalized federated learning (PFL) strategies aim to accomplish this through two primary approaches: (1) globally anchored personalization, where local models leverage knowledge from a central global model; and (2) peer-based personalization, where local models learn from similar clients identified at the server level. Globally anchored strategies are appropriate in applications where we expect each client to have limited dissimilarity with any other client in the network, such as in next-word prediction on mobile devices. However, in applications such as recommender systems, distributional differences between two given clients can be very large. Peer-based methods address this issue by attempting to only learn from clients that are similar. One approach to peer-based personalization is to cluster clients, prior to training or concurrently. The cluster model of statistical heterogeneity does however fail to capture more complex overlaps between clients. As an example, in a music recommendation application, suppose one user (client) is only interested in jazz and classical, another in jazz and funk, and yet another in funk and rock. Clearly, no hard clustering of these users will capture these relationships accurately. Other approaches to peer-based personalization allow for more complex relationships between the local models of clients, but to the best of our knowledge no prior work incorporates distributional similarity measures in a principled manner.
@@ 1.1 Contributions
The main contributions of this work are threefold:
  • We propose a peer-based PFL strategy that learns personalized models that are centerally constrained to have limited pairwise dissimilarity. These constraints are derived in a principled, fully data-driven manner using approximations of the Wasserstein distances between empirical client distributions. This design is grounded in theoretical insights developed in this work.
  • We develop an efficient inexact projected stochastic gradient method tailored to our framework. It employs variance reduction to deal with partial client participation in each communication round. Due to this variance reduction, we are able to show convergence for smooth, possibly non-convex losses at rate $O(1/K)$ to an $\epsilon$-stationary set of personalized models, where the error $\epsilon$ arises only due to stochastic gradients from the clients and inexact projection at the server level.
  • We evaluate our approach on both synthetic and real-world heterogeneous datasets across regression and classification tasks. Our method is compared against a variety of standard federated and personalized FL baselines, demonstrating several distinctive advantages.
@@ 1.2 Outline
In §2 we describe some of the related work, relating our contribution to existing literature. Then, in §3, we introduce our proposed strategy for personalized federated learning, which is followed by a convergence analysis in §4. Numerical experiments validate our proposed method in §5. We conclude the paper by summarizing and discussing our findings in §6.
@@ 2 Related work
Statistical heterogeneity presents significant challenges to both traditional and distributed learning systems, affecting inference and generalization. It arises in various applications where data can be assumed to be drawn from distinct distributions. For example, in meta-analysis, studies with participants from different populations are aggregated (Melsen et al.,, 2014); in medicine, different clinical features in patients affect individual health outcomes (Machado et al.,, 2011); and in finance, markets adopt different regimes over time (Ang and Timmermann,, 2012). In the context of FL, the statistical heterogeneity among clients is generally considered one of the defining aspects of the field. Zhao et al., (2018) and Karimireddy et al., (2020) show that both convergence and generalization of conventional FL algorithms are severely affected by statistical heterogeneity. The heterogeneity introduces difficulties in optimizing the global model, as local updates cause models to drift, reducing the effectiveness of aggregation methods and slowing down convergence. The former work shows that the 1-Wasserstein distance between client distributions can be used to quantify this effect.
One approach to address the challenges caused by statistical heterogeneity is stratification, whereby different models are fit for every value of some categorical parameter, which are assumed to correspond to the different underlying distributions. Stratified models allow each subgroup to be modeled independently, capturing the unique characteristics of each distribution and reducing the bias that arise from pooling heterogeneous data. In clinical trials, it is common to do this by grouping participants into different strata based on relevant clinical features, and fitting a model for each strata (Kernan et al.,, 1999). Stratification does, however, reduce the number of observations for each model, limiting the viability of naive stratification in many settings. PFL models, which adapt to individual clients rather than enforce a single global model, are a special case of stratified models where the stratification is performed at the level of individual clients.
@@
A more flexible personalization strategy is to identify clients that are similar and borrow strength from them directly rather than from a single aggregated model. One way to do this is by clustering the clients. The PFL method proposed by Ghosh et al., (2020) does this iteratively by alternating between optimizing the cluster models and clustering the clients, which is done by finding the cluster whose model achieves the smallest local loss. Regularized methods for peer-based PFL include convex clustering via a group-lasso penalty on the differences of models (Armacki et al.,, 2022) and Laplacian regularization (Dinh et al.,, 2022; Hashemi et al.,, 2024). The empirical results of Dinh et al., show that in fact Laplacian regularization with randomized weights outperform globally anchored personalization strategies on many tasks.
@@ 3 Proposed method
In this section, we introduce Karula [^1] The name is a portmanteau of Kantorovich, Rubenstein and Laplace., a strategy for federated learning of personalized models where model differences are constrained by the distributional similarity of client data. This reinforces the intuitive notion that clients with similar data should learn similar models. We also propose an efficient projected stochastic gradient method with variance reduction to handle partial client participation during communication rounds. This method is also suitable for use in other FL strategies.
@@ 3.1 Main idea
Consider an FL setting with $N$ clients, each able to communicate with a central server. To each client $i\in\{1,\dots,n\}$, we associate an underlying distribution $\mu_{i}\colon 2^{\mathcal{X}}\to{\mbox{\bf R}}_{+}$, where $\mathcal{X}$ is the feature space. Each client holds a dataset $\mathcal{D}_{i}$ of $N_{i}$ independent samples drawn from $\mu_{i}$. The learning goal is to find a personalized model $\theta_{i}\in\Theta$ for each client that minimizes their expected loss $f_{i}(\theta_{i})={\mathbbmss{E}}_{\xi\sim\mu_{i}}[\ell(\theta_{i},\xi)]$, where $\ell\colon\Theta\times\mathcal{X}\to{\mbox{\bf R}}$ is a loss function, and $\Theta$ is the parameter space. However, due to limited available data on each client, attempting to directly minimize $f_{i}(\theta_{i})$ leads to poor generalization. To address this challenge, we introduce a regularization strategy that couples the learning tasks across clients by approximately solving a constrained optimization problem:
@@
Here $t\geq 0$ is a regularization parameter, each $\alpha_{i}$ is a positive weight, and each $D_{ij}\geq 0$ is a dissimilarity parameter which quantifies the dissimilarity between clients $i$ and $j$ that is calculated before solving the minimization problem. This formulation generalizes the standard FL formulation, since by selecting $t=0$ the problem is equivalent to choosing a single global model that minimizes the average loss. In the other extreme, when $t\to\infty$, we obtain completely local models except for clients that are deemed to have identical distributions, i.e., when $D_{ij}=0$. In the sequel, we show how to choose the similarity parameters in a completely data-driven manner, based on a surrogate for the Wasserstein distance.
@@ 3.2 Model dissimilarity and Wasserstein distance
To quantify the relationship between data distributions and the corresponding learned models, we consider the 1-Wasserstein distance—also known as the earth mover’s distance or the Kantorovich–Rubinstein distance. Suppose the data space $\mathcal{X}$ is equipped with a metric $d_{\mathcal{X}}$. The 1-Wasserstein distance between two distributions $\mu_{i}$ and $\mu_{j}$ is defined as
@@
where $\Pi(\mu_{i},\mu_{j})$ is the set of all probability measures with marginals $\mu_{i}$ and $\mu_{j}$. This distance is a proper metric on the space of distributions on $\mathcal{X}$, satisfying symmetry, positive definiteness, and the triangle inequality (Villani,, 2008).
@@
The latter assumption holds naturally for some loss functions, but for others, e.g. linear regression with quadratic loss, we need to require $\Theta$ to be bounded.
@@
A straight-forward approach to shrink the model dissimilarities would be to learn completely local models first, calculate their pairwise distances, and then use them as the dissimilarity parameters. So why use a less direct approach with Wasserstein distances? In the FL setting, it is common for clients to have more features than observations. For instance, in a linear regression task with $d$ important features and $N$ observations per client, it is well known that the minimax squared estimation error is $O(d/N)$ (Bickel et al.,, 2009; Raskutti et al.,, 2011). When $N<d$, no meaningful upper bound on the estimation error can be guaranteed without additional structural assumptions. The situation worsens for overparameterized models, where the number of parameters exceeds the number of features. In contrast, the Wasserstein distance between two empirical distributions has a rate of $O(N^{-1/d})$ (Wang et al.,, 2021, Example 1). While this rate is very slow in high dimensions, it clearly performs much better than $O(d/N)$ when $N$ is on a similar scale as $d$, or is smaller. We show empirically in §5.1 how our approach is better at estimating the true dissimilarities.
@@
In supervised learning, a natural way to define the data metric $d_{\mathcal{X}}$ is to concatenate the feature vector and response, and apply a standard norm—such as the Euclidean distance—on the joint space. An alternative approach, proposed by Elhussein and Gursoy, (2024), combines cosine similarity for the features with the Hellinger distance for the responses. They argue that cosine-similarity is more informative in high-dimensional feature spaces due to the concentration of measures phenomenon, and that the Hellinger distance compares densities within their statistical manifold, potentially yielding more meaningful distinctions between distributions.
@@
With the present motivation in mind, it is natural to set the dissimilarity parameters proportional to the Wasserstein distance between the empirical distributions, $W_{1}(\hat{\mu}_{i},\hat{\mu}_{j})$, where $\hat{\mu}_{i}$ is the empirical distribution of client $i$th. While Rakotomamonjy et al., (2023) propose an iterative method to approximate these distances in a federated manner, their approach requires computing $n(n-1)/2$ optimal transport plans per iteration—an operation that is computation and communication-intensive.
To reduce this cost, we propose to approximate the Wasserstein distances via linear embedding of the data sets, following prior work such as Kolouri et al., (2020) and Liu et al., (2025). Specifically, we introduce a reference data set of $N_{0}$ independent samples $\mathcal{D}_{0}=\{\xi_{1}^{0},\dots,\xi_{N_{0}}^{0}\}$ drawn from a reference distribution $\mu_{0}$. The optimal transport plan between the empirical distributions $\hat{\mu}_{0}$ and $\hat{\mu}_{i}$ is then
@@
By approximating the Monge map via $M_{i}=N_{0}\pi_{i}^{\star}\mathcal{D}_{i}$, we can define an embedding $\Phi$ by $\Phi(\mathcal{D}_{i})=(M_{i}-\mathcal{D}_{0})/\sqrt{N_{0}}$ that has the property that $\|\Phi(\mathcal{D}_{i})-\Phi(\mathcal{D}_{j})\|_{1}$ defines a metric on the space of data sets, and indeed approximates the Wasserstein distance $W_{1}(\hat{\mu}_{i},\hat{\mu}_{j})$. We thus choose the dissimilarity parameters as $D_{ij}=\|\Phi(\mathcal{D}_{i})-\Phi(\mathcal{D}_{j})\|_{1}$. By doing this, each client only has to compute a single optimal transport plan. Also, on the communication side, only a reference data set and the embeddings need to be transmitted over the network once, totaling $O(N_{0}dn)$ communication.